# Characterizing Open-Ended Evolution Through Undecidability Mechanisms in Random Boolean Networks

### v1

@@ Methods
While canonical GRNs are classical, quantum-inspired rules can model contextuality and “superposition-like” uncertainty during decision points (e.g., fate priming) (cao2020quantum). Moreover, quantum coherence and related effects have been directly observed in biological light-harvesting complexes (e.g., the Fenna-Matthews-Olson bacteriochlorophyll complex), suggesting that non-classical information processing can appear in biology under some conditions—even if not in gene regulation per se (engel2007evidence; panitchayangkoon2010long). The point here is not that GRNs are quantum, but that quantum-style operators offer a principled way to represent context-dependent mixture/branching and long-range correlations that classical gates do not capture natively (franco2021random).
In our work, we extend BN with (i) a superposition token and (ii) an entanglement pairing $E\subseteq V\times V$. LUT entries then take values in $\{0,1,\text{superposed}\}$. At each update, the evaluator $\kappa$ maps the superposed state to a Bernoulli($1/2$) collapse in $\{0,1\}$. If $i$ is entangled with $j$, then $x_{i}(t+1)=x_{j}(t)$ (a classical proxy for instantaneous correlation). Formally, $x_{i}(t+1)=\kappa\left(f_{i}\left(x_{i}(t)\right)\right)$ if $(i,j)\notin E$; otherwise, $x_{i}(t+1)=x_{j}(t)$. In our code, we uniformly assign a value (0 or 1) to the superposed state and, for entangled nodes, we simply copy the partner’s current bit. This mirrors the behavioral signatures emphasized by franco2021random, i.e., superposition-like sampling and long-range correlations without changing the underlying Boolean state space. Now we can proceed to introduce our open-endedness metric.
@@ A Novel Metric for Quantifying Open-Endedness in Any Discrete State Space
In the deterministic case, regardless of their initial configuration, all Boolean networks fall into an attractor (gershenson2004introduction). Attractors are stable sets of states that represent the long-term behavior of the network. Once a Boolean network reaches an attractor, it will remain within that set of states indefinitely unless perturbed. Attractors can be single states (point attractors) or cycles of multiple states (cycle attractors) and are often linked to biological phenotypes, such as cell types in GRNs (Kauffman1969). For any attractor in a Boolean network, we can calculate the number of unique states the network cycles through repeatedly (cycle-length) and the number of time steps the network remains trapped within the attractor (residence-time). In this way, given a RBN with $N$ nodes where $m$ different attractors appear during $T$ time steps, we define open-endedness as
@@
where $k_{j}$ and $d_{j}$ are respectively the cycle-length and residence-time of the attractor $A_{j}$ discovered within $T$ time-steps. In layman’s terms, open-endedness is the residence-time-weighted temporal average of the individual cycle-lengths for each attractor discovered in the current dynamics, as the number of time steps tends to infinity. In the deterministic case, when $m=1$, we have a single attractor of length $k_{1}\in\{1,2,\dots,2^{N}\}$. If we enter this attractor at time step $t_{0}$, then
@@
In other words, a process that constantly produces new states but never dwells on any of them is classified as non-open-ended by our metric: its novelty is pure noise. The same conclusion holds for a “slow” or delayed random walk on a finite state space that includes self-loops, as long as the expected residence-time in any configuration remains bounded while $T$ grows. In that case, the numerator $\sum_{j}d_{j}k_{j}$ can grow at most linearly with $T$, so $\Omega(T)=\mathcal{O}(1/T)$ and still vanishes in the limit. Non-zero $\Omega$ therefore requires not only the continual appearance of distinct attractors, but also that the total time spent in structured, recurrent dynamical patterns grows faster than linearly with $T$. Now let us assume that the dynamics are no longer deterministic or completely random. By construction, there exist an ordered sequence of time steps, $\{t_{1}^{i},t_{1}^{f},t_{2}^{i},t_{2}^{f},\dots,t_{m}^{i},t_{m}^{f}\}$, such that $A_{j}$ appears in $t_{j}^{i},t_{j}^{f}]~~~\forall j\in\{1,\dots,m\}$. This implies the following chain of equalities and inequalities:
@@ Results
We evaluated open-endedness ($\Omega$) across classical Random Boolean Networks (RBNs) and four non-Boolean extensions (Paraconsistent, Modal, Quantum-inspired, and Annealed Rule Mutation (ARM)) together with Probabilistic Boolean Networks (PBNs). The analysis pipeline proceeded as follows.
@@
    First, for each average connectivity $K$ on a fixed grid ($K\in[1.1,4.5]$ in steps of $0.2$), we generated $1,000$ independent RBNs of $N=100$ nodes with random initial states and simulated each network for $T=10^{6}$ time steps. The full parameter sweeps used to select the settings reported below, as well as sensitivity checks, are provided in the Supplementary Material.
@@
    Finally, to compare logics concisely in the main text we plot only these best curves for the homogeneous and heterogeneous cases, respectively; all intermediate sweeps and additional diagnostics appear in the Supplementary Material.
@@
We distinguish two architectural regimes. In the homogeneous case, networks have structural, temporal, and functional homogeneity; in practice we use a Poisson in-degree distribution (fixed mean $K$), synchronous updating, and a constant bias of $0.5$. In the heterogeneous case, we introduce structural and temporal heterogeneity; in practice we use an Exponential in-degree distribution (same mean $K$) and a non-deterministic asynchronous update scheme. In both regimes, the classical baseline is a deterministic BN; PBNs add context switching across deterministic constituents; ARM flips LUT outputs at read-time with probability $\mu$; Paraconsistent logic injects locally contradictory table entries with probability $c$; Modal logic rewrites entries to modal tokens according to $(a,p_{p},p_{n})$ (accessibility degree and the probabilities of possible/necessary); and Quantum-inspired logic introduces superposed outputs with probability $sp$ and pairs up to $e$ entangled nodes (see Methods).
@@
Figures 1 and 2 report, for each regime, the single best curve per logic under the selection rule above. Legends list the winning parameter settings for each mechanism; for consistency we abbreviate parameters as: “ctx” = number of PBN contexts, $\sigma$ = PBN switching probability per epoch, $\mu$ = ARM mutation probability, $c$ = paraconsistent contradiction probability, $a$ = modal accessibility degree, $p_{p}$$p_{n}$ = modal probabilities of writing possible/necessary, $sp$ = superposition probability, and $e$ = number of entangled pairs. Axes are shared across panels: the horizontal axis is average connectivity $K$; the vertical axis is our open-endedness score $\Omega$ (see Methods). Because our goal in this section is to document the outputs and the evaluation process, we defer interpretation to the Discussion section, where we will start by exploring how our metric is related to previous approaches, showing how it is a complementary measure of novelty.
@@ Discussion
Our starting point is the formal program that sought system-independent definitions of “unbounded evolution” (UE) and “innovation” in discrete dynamical systems (adams2017formal). That work emphasizes unboundedness and genuinely novel state production as the core signatures of OEE, abstracted away from any particular substrate or encoding. Our metric takes that spirit and turns it into a computable diagnostic that is applicable to any discrete state space: by counting distinct phenotypic outcomes (attractors or coarse-grained phenotypes) produced across a sliding temporal window, it operationalizes “ongoing innovation” without assuming access to asymptotic time. In particular, once the number $m$ of distinct attractors encountered by the dynamics diverges, our open-endedness score $\Omega$ diverges as well, matching the intuition that an unbounded supply of phenotypes corresponds to open-endedness, while remaining usable at finite horizons. This finite-horizon property complements adams2017formal asymptotic criteria for UE/innovation and allows us to probe parameter regimes where novelty generation accelerates before any formal proof of unboundedness is in sight.
@@
Compared to the Evolutionary Activity Statistics (EAS) tradition and the Bedau-Packard classification of long-term dynamics (bedau1997comparison; bedau1998classification), our approach dispenses with component-level activity counters and the accompanying choices of neutral baselines. The Bedau program classifies systems via diversity and activity time series, a lens that has been productively used but depends on how components and activity are defined and normalized across domains. Our metric instead interrogates the reachable state space itself—counting distinct phenotypic outcomes realized within a moving window—and therefore avoids reliance on a priori component partitions or domain-specific neutral models while still distinguishing regimes with sustained novelty from those that saturate. In this way it is compatible with, but orthogonal to, EAS-style summaries.
@@
The MODES toolbox reframed measurement around four high-level “hallmarks” (change, novelty, complexity, and ecological potential) offering portable algorithms and inviting lower-level mechanistic metrics to complement them (dolson2019modes). Our metric precisely fills that niche: it is a low-level, substrate-agnostic counter of realized phenotypic novelty that can be dropped into any discrete-state model (including RBN variants) and read alongside MODES’ hallmarks. In particular, MODES highlights continual novelty as a central hallmark; our sliding-window $\Omega$ makes this hallmark numerically explicit in systems where “phenotype” is tied to attractors or coarse-grained dynamical outcomes, and it can be summarized jointly with MODES’ other hallmarks to separate sustained novelty from transient bursts.
@@
Our emphasis on ongoing novelty also dovetails with the algorithmic-information view that relates strong forms of OEE to undecidability. hernandez2018undecidability showed that, for a broad class of sophistication-like complexity measures, decidability imposes absolute limits on stable complexity growth and that strong OEE entails undecidability/irreducibility of adapted states. While our $\Omega$ deliberately does not attempt to estimate sophistication (kolmogorov1965three) or logical depth (bennett1988logical), it is consistent with that result: in regimes where undecidable structure drives unbounded production of phenotypic outcomes, $\Omega$ will continue to rise with the number of distinct attractors realized; in decidable regimes, $\Omega$ saturates. Thus, our metric provides a computable proxy for detecting the phenomenology that undecidability results lead us to expect, without invoking heavy algorithmic-complexity estimators whose incapability of distinguishing novelty from noise has been shown (posobin2018random).
@@
Furthermore, our metric is squarely aimed at the York workshop’s category of “ongoing generation of adaptive novelty” (taylor2016open). It implements exactly what that category asks us to detect: sustained production of new (and potentially adaptive) phenotypes over time, operationalized in finite windows and applicable across substrates. We explicitly do not promote York’s second category (“ongoing growth of complexity”) as a necessary lens on OEE. Beyond long-standing conceptual worries about what counts as “complexity,” there is little biological support for the claim that natural selection relentlessly drives increases in organismal complexity; recent arguments review both theory and evidence indicating that complexity increases are contingent, often nonadaptive, and concentrated in a few lineages, with many trends pointing in the opposite direction (lynch2025complexity). Our focus on novelty avoids conflating open-endedness with monotone complexity growth, in line with these cautions.

### v2

@@ Methods
While canonical GRNs are classical, quantum-inspired rules can model contextuality and “superposition-like” uncertainty during decision points (e.g., fate priming) (Cao et al., 2020). Moreover, quantum coherence and related effects have been directly observed in biological light-harvesting complexes (e.g., the Fenna-Matthews-Olson bacteriochlorophyll complex), suggesting that non-classical information processing can appear in biology under some conditions—even if not in gene regulation per se (Engel et al., 2007; Panitchayangkoon et al., 2010). The point here is not that GRNs are quantum, but that quantum-style operators offer a principled way to represent context-dependent mixture/branching and long-range correlations that classical gates do not capture natively (Franco et al., 2021). We emphasize that our goal is not to implement formal quantum logic (e.g., non-distributive propositional lattices (Hartonas, 2018)) nor quantum mechanics (e.g., unitary evolution in a Hilbert space (Das et al., 2021)).
Instead, we introduce a quantum-inspired mechanism that captures two qualitative motifs often associated with “quantum-like” behavior in complex systems: (i) stochastic resolution of an underspecified update (a superposition-inspired token that collapses to a Boolean value) and (ii) long-range instantaneous correlation implemented as a paired-state coupling between nodes. Concretely, LUT entries may take values in $\{0,1,\texttt{superposed}\}$. At update time, an evaluator $\kappa$ maps superposed to a Bernoulli(1/2)(1/2) draw in $\{0,1\}$, so the realized network state remains Boolean. Separately, we sample a set of paired couplings $E\subseteq V\times V$ (a set of node pairs). If node $i$ is paired with node $j$, then $x_{i}(t+1)=x_{j}(t)$, which acts as a classical proxy for instantaneous correlation. This construction is therefore best understood as a logic-inspired correlated-branching update rule, not as a claim about genuine quantum dynamics. Now we can proceed to introduce our open-endedness metric.
@@ A Novel Metric for Quantifying Open-Endedness in Any Discrete State Space
In the deterministic case, regardless of their initial configuration, all Boolean networks fall into an attractor (Gershenson, 2004a). Attractors are stable sets of states that represent the long-term behavior of the network. Once a Boolean network reaches an attractor, it will remain within that set of states indefinitely unless perturbed. Attractors can be single states (point attractors) or cycles of multiple states (cycle attractors) and are often linked to biological phenotypes, such as cell types in GRNs (Kauffman, 1969). For any attractor in a Boolean network, we can calculate the number of unique states the network cycles through repeatedly (cycle-length) and the number of time steps the network remains trapped within the attractor (residence-time). In this way, given a RBN with $N$ nodes where $m$ different attractors appear during $T$ time steps, we define open-endedness as
@@
where $k_{j}$ and $d_{j}$ are respectively the cycle-length and residence-time of the attractor $A_{j}$ discovered within $T$ time-steps. In all experiments of the present manuscript we use an operational finite-horizon recurrence detector that extracts attractor episodes from a single simulated trajectory, rather than attempting a full invariant (at $T\to\infty$) decomposition of the underlying dynamics. Although several mechanisms introduce non-Boolean rule-level tokens (e.g., modal possible/necessary, quantum-inspired superposed, etc.), the network state vector that we record at each time step remains Boolean. Rule outputs are evaluated through a semantics map (e.g., modal evaluation, superposition “collapse”, etc.) that produces a Boolean update. Thus, recurrence is always detected on realized Boolean microstates, not on intermediate symbolic labels.
Let $\{x(t)\}_{t=0}^{T}$ be the realized Boolean trajectory. We maintain a hash map $\mathrm{firstSeen}[x]$ that stores the earliest time index at which each state $x$ appeared. At time $t$ we declare a cycle discovery when $x(t)$ matches a previously seen state $x(\tau)$ with $\tau<t$ and we are not currently inside an episode. We then set the cycle-length $k:=t-\tau$ and start a new episode anchored at $\tau$. Once an episode starts, its dwell time $d$ is the number of subsequent time steps the trajectory continues to spend inside the previously explored recurrent region. Operationally, $d$ increments by $1$ each step until the trajectory hits a never-before-seen state (a state not in $\mathrm{firstSeen}$), which we treat as an escape. While inside an episode, additional returns to already visited states do not create new episodes. In particular, if the trajectory traverses multiple short loops that share states, or returns to a prior loop after visiting another loop (a “figure-eight” / chained-cycle pattern), this is counted as one continuous recurrent episode rather than multiple independent attractors.
In this way, a new episode can begin only after at least one genuinely novel state appears. Below is the minimal logic used to extract $(k_{j},d_{j})$ from a trajectory (notations correspond to the codebase):
firstSeen = {}; episode = None
for t = 0..T-1:
s = hash(x(t))
if episode is None and s in firstSeen:
tau = firstSeen[s]; k[tau] = t - tau; d[tau] = 1; episode = tau
else if episode is not None:
d[episode] += 1
if s not in firstSeen: episode = None
if s not in firstSeen: firstSeen[s] = t
We report the finite-horizon proxy $\Omega(T)=\frac{1}{T^{2}}\sum_{j}d_{j}k_{j}$ computed from the above episode detector. For deterministic systems this coincides with standard cycle detection once the orbit enters a true limit cycle. For stochastic systems, $\Omega(T)$ instead summarizes empirical recurrence structure within a finite horizon, i.e., how often the dynamics returns to previously visited microstates and how long it continues to revisit the explored recurrent region. This is a recurrence-based diagnostic, not a strict Markov-chain attractor decomposition. In layman’s terms, open-endedness is the residence-time-weighted temporal average of the individual cycle-lengths for each attractor discovered in the current dynamics, as the number of time steps tends to infinity. In the deterministic case, when $m=1$, we have a single attractor of length $k_{1}\in\{1,2,\dots,2^{N}\}$. If we enter this attractor at time step $t_{0}$, then
@@
In other words, a process that constantly produces new states but never dwells on any of them is classified as non-open-ended by our metric: its novelty is pure noise. The same conclusion holds for a “slow” or delayed random walk on a finite state space that includes self-loops, as long as the expected residence-time in any configuration remains bounded while $T$ grows. In that case, the numerator $\sum_{j}d_{j}k_{j}$ can grow at most linearly with $T$, so $\Omega(T)=\mathcal{O}(1/T)$ and still vanishes in the limit. For a deterministic map on a finite state space, every trajectory eventually becomes periodic. Let $\mu$ be the first time at which the trajectory revisits a previous state (the entry time into the eventual cycle), and let $\lambda$ be the corresponding cycle length. For any horizon $T>\mu$, the trajectory spends the remaining $(T-\mu)$ steps on the same cycle, so the operational sum in Eq. (2) reduces to a single attractor episode with $k=\lambda$ and $d(T)=T-\mu$, yielding
$\Omega(T)=\frac{1}{T^{2}}\,\lambda\,(T-\mu)=\frac{\lambda}{T}\left(1-\frac{\mu}{T}\right)\;=\;\mathcal{O}\!\left(\frac{1}{T}\right).$ (5)
Thus, once the eventual cycle has been reached and identified, $\Omega(T)$ must decay approximately as $1/T$ as the observation window grows. More generally, for any fixed finite state space one has $\sum_{j}d_{j}k_{j}\leq(2^{N})\,T$, hence $\Omega(T)\leq 2^{N}/T$ and the strict $T\to\infty$ limit vanishes even for stochastic dynamics; the practical role of $\Omega(T)$ is therefore as a finite-horizon recurrence diagnostic. In the Supplementary Material (Fig. S22) we explicitly check the horizon dependence for both the deterministic RBN baseline and a stochastic PBN mechanism, and include a small-$N$ control to illustrate when the asymptotic $1/T$ regime becomes observable.
To address external validation, index comparisons, and robustness to trivial relabelings, we evaluated $\Omega$ on (i) a small panel of Elementary Cellular Automata (ECA) rules spanning fixed/periodic/complex regimes, and (ii) curated Boolean gene-regulatory network (GRN) models. In addition, we computed three common trajectory-level proxies on the same runs (unique-state fraction, mean node entropy, and a compression proxy) and recomputed $\Omega$ under the complement quotient that identifies each Boolean state $s$ with its complement 1s1-s. These results are reported in Supplementary Figs. S23–S25.
Taken together, these results show that non-zero $\Omega$ requires not only the continual appearance of distinct attractors, but also that the total time spent in structured, recurrent dynamical patterns grows faster than linearly with $T$. Now let us assume that the dynamics are no longer deterministic or completely random. By construction, there exists an ordered sequence of time steps, $\{t_{1}^{i},t_{1}^{f},t_{2}^{i},t_{2}^{f},\dots,t_{m}^{i},t_{m}^{f}\}$, such that $A_{j}$ appears in $t_{j}^{i},t_{j}^{f}]~~~\forall j\in\{1,\dots,m\}$. This implies the following chain of equalities and inequalities:
@@ Results
We evaluated open-endedness ($\Omega$) across classical Random Boolean Networks (RBNs) and four non-Boolean extensions (Paraconsistent, Modal, Quantum-inspired, and Annealed Rule Mutation (ARM)) together with Probabilistic Boolean Networks (PBNs). All of these mechanisms make the dynamics stochastic. In these cases, we do not claim to compute invariant attractors in the strict dynamical-systems sense. Instead, we measure empirical recurrence episodes within a finite horizon using the operational detector described in Methods. Consequently, $\Omega(T)$ should be interpreted as a finite-time recurrence-based signature of sustained novelty, not as a canonical attractor decomposition of the underlying Markov process. Because the state space scales as $2^{N}$, very long cycles (or extremely slow recurrences) may not be observed within $T=10^{6}$ steps at $N=100$. Our reported $\Omega(T)$ is therefore a finite-horizon proxy that captures recurrence structure that is actually realized within the observation window. Supplementary Figs. S1–S2 report horizon sensitivity checks. The analysis pipeline proceeded as follows.
@@
    First, for each average connectivity $K$ on a fixed grid ($K\in[1.1,4.5]$ in steps of $0.2$), we generated $1,000$ independent RBNs of $N=100$ nodes with random initial states and simulated each network for $T=10^{6}$ time steps. The full parameter sweeps used to select the settings reported below, as well as sensitivity checks, are provided in the Supplementary Material (Figures S1–S25 and Tables S1–S2).
@@
    Finally, to compare logics concisely in the main text we plot only these best curves for the homogeneous and heterogeneous cases, respectively; all intermediate sweeps and additional diagnostics appear in the Supplementary Material (Figures S3-S21 and Tables S1-S2).
@@
We distinguish two architectural regimes. In the homogeneous case, networks have structural, temporal, and functional homogeneity; in practice we use a Poisson in-degree distribution (fixed mean $K$), synchronous updating, and a constant bias of $0.5$. In the heterogeneous case, we introduce structural and temporal heterogeneity; in practice we use an Exponential in-degree distribution (same mean $K$) and a non-deterministic asynchronous update scheme. In both regimes, the classical baseline is a deterministic BN; PBNs add context switching across deterministic constituents; ARM flips LUT outputs at read-time with probability $\mu$; Paraconsistent logic injects locally contradictory table entries with probability $c$; Modal logic rewrites entries to modal tokens according to $(a,p_{p},p_{n})$ (accessibility degree and the probabilities of possible/necessary); and quantum-inspired mechanisms introduce superposed outputs with probability $sp$ and pairs up to $e$ pair-coupled nodes (see Methods).
@@
Figures 1 and 2 report, for each regime, the single best curve per logic under the selection rule above. Legends list the winning parameter settings for each mechanism; for consistency we abbreviate parameters as: “ctx” = number of PBN contexts, $\sigma$ = PBN switching probability per epoch, $\mu$ = ARM mutation probability, $c$ = paraconsistent contradiction probability, $a$ = modal accessibility degree, $p_{p}$$p_{n}$ = modal probabilities of writing possible/necessary, $sp$ = superposition probability, and $e$ = number of pair-coupled pairs. Axes are shared across panels: the horizontal axis is average connectivity $K$; the vertical axis is our open-endedness score $\Omega$ (see Methods). Now that we have documented the results obtained, let’s interpret them.
In homogeneous networks, the PBN transformation achieves the largest $\Omega$ at low-medium connectivities, whereas Modal logic dominates at high connectivities (Fig. 1). The PBN outcome is consistent with the mechanism: piecewise deterministic motion inside a context, punctuated by rare switches, tends to string together long residence times in distinct attractor families, inflating the residence-time-weighted cycle contribution in $\Omega$. Biologically, this mirrors bet-hedging via phenotype switching under resource-scarce conditions: exactly where sparse coupling (low $K$) limits within-context exploration, so context changes do the heavy lifting (Acar et al., 2008). By contrast, when coupling is dense (large $K$), Modal gating becomes advantageous. “Necessary” constraints act as context-dependent canalizing conditions that suppress runaway sensitivity, while “possible” gates open alternative routes when local context permits; together they tame chaotic overshoot and create metastable corridors that repeatedly access distinct limit cycles. As a practical rule: in synchronous, homogeneous systems with limited connectivity, stochastic program switching (PBN) yields the highest open-endedness; as connectivity grows, context-dependent gating (Modal) overtakes by structuring exploration without freezing it.
Moving to heterogeneous networks, the picture changes qualitatively (Fig. 2). The PBN curve collapses toward zero over the entire $K$ range, while Modal again excels at higher $K$, and Paraconsistent and Quantum-inspired mechanisms show localized gains at low-intermediate $K$. This collapse is not due to a change in the number of point attractors, those are determined by the rule tables and are independent of the updating scheme (Gershenson, 2004b), but rather to how our stochastic asynchronous updates interact with structural heterogeneity. In our simulations we use CUBEWALKERS’ asynchronous_set update scheme, which randomly selects a set of nodes to be updated at each time step, so that the dynamics along a single trajectory are effectively noisy and long synchronous cycles are rarely realized. Combined with an Exponential in-degree distribution, this drives most initial conditions quickly into small recurrent classes (often fixed points or very short cycles) that have large basins of attraction.
Within each PBN context, trajectories therefore spend almost all their time in short attractors with cycle lengths $k_{j}$ of order one; context switches mostly move the system between the basins of these same few attractors, rather than uncovering new long cycles. As a result, the summed residence-cycle contribution $\sum_{j}d_{j}k_{j}$ grows at most linearly with $T$, so $\Omega\rightarrow 0$ despite ongoing context changes. Modal logic is the exception because its necessity/possibility tests provide conditional activation that intermittently re-opens pathways even under asynchronous settling; hence its large spikes at high $K$. Paraconsistent logic yields a broad shoulder around $K\approx 2.3$: introducing controlled local contradictions near the order-critical boundary sustains exploratory dynamics without exploding to noise, in line with the known criticality-extending effects of heterogeneity (Sánchez-Puig et al., 2023; López-Díaz et al., 2023). Quantum-inspired rules perform best at very low $K$: superposition tokens plus a few long-range paired-state couplings furnish correlated branching when wiring is sparse, boosting early-time exploration before asynchronous contraction dominates. The strategy that follows is therefore different from the homogeneous case: under heterogeneity/asynchrony, prefer Modal gating at high connectivity and use Paraconsistent or Quantum-inspired rules to lift $\Omega$ in sparse networks.
Some patterns are trivial yet worth stating explicitly. In both regimes, the Classical curve sinks over the entire $K$ range because a deterministic network visiting a single attractor contributes $\Omega\rightarrow 0$ for large enough values of $T$ (see Supplementary Material, Figures S1–S2); $\Omega$ rewards persistent traversal of multiple cyclic phenotypes, not mere residence in one. The consistently poor performance of ARM is also instructive: per-read truth-table mutation anneals away stable local structure, pushing dynamics toward effectively memoryless transitions; this drives down the product of cycle length and dwell time and is qualitatively consistent with the classic annealed-approximation intuition that high flip rates erase organized attractor geometry (Derrida and Pomeau, 1986). In short, ARM provides plasticity but not the kind of structured, state-dependent plasticity that $\Omega$ rewards.
In the next section, we will explore how our metric is related to previous approaches, showing how it is a complementary measure of novelty. As we shall see, our approach focuses specifically on discrete state spaces, leaving aside systems with continuous (or hybrid) state spaces. Thus, additionally, we sketch a principled extension of our metric that preserves its core properties while making it applicable beyond purely discrete models.
@@ Discussion
Our starting point is the formal program that sought system-independent definitions of “unbounded evolution” (UE) and “innovation” in discrete dynamical systems (Adams et al., 2017). That work emphasizes unboundedness and genuinely novel state production as the core signatures of OEE, abstracted away from any particular substrate or encoding. Our metric takes that spirit and turns it into a computable diagnostic that is applicable to any discrete state space: by counting distinct phenotypic outcomes (attractors or coarse-grained phenotypes) produced across a sliding temporal window, it operationalizes “ongoing innovation” without assuming access to asymptotic time. In particular, once the number $m$ of distinct attractors encountered by the dynamics diverges, our open-endedness score $\Omega$ diverges as well, matching the intuition that an unbounded supply of phenotypes corresponds to open-endedness, while remaining usable at finite horizons. This finite-horizon property complements Adams et al. (2017) asymptotic criteria for unbounded evolution/innovation and allows us to probe parameter regimes where novelty generation accelerates before any formal proof of unboundedness is in sight.
@@
Compared to the Evolutionary Activity Statistics (EAS) tradition and the Bedau-Packard classification of long-term dynamics (Bedau et al., 1997, 1998), our approach dispenses with component-level activity counters and the accompanying choices of neutral baselines. The Bedau program classifies systems via diversity and activity time series, a lens that has been productively used but depends on how components and activity are defined and normalized across domains. Our metric instead interrogates the reachable state space itself—counting distinct phenotypic outcomes realized within a moving window—and therefore avoids reliance on a priori component partitions or domain-specific neutral models while still distinguishing regimes with sustained novelty from those that saturate. In this way it is compatible with, but orthogonal to, EAS-style summaries.
@@
The MODES toolbox reframed measurement around four high-level “hallmarks” (change, novelty, complexity, and ecological potential) offering portable algorithms and inviting lower-level mechanistic metrics to complement them (Dolson et al., 2019). Our metric precisely fills that niche: it is a low-level, substrate-agnostic counter of realized phenotypic novelty that can be dropped into any discrete-state model (including RBN variants) and read alongside MODES’ hallmarks. In particular, MODES highlights continual novelty as a central hallmark; our sliding-window $\Omega$ makes this hallmark numerically explicit in systems where “phenotype” is tied to attractors or coarse-grained dynamical outcomes, and it can be summarized jointly with MODES’ other hallmarks to separate sustained novelty from transient bursts.
@@
Our emphasis on ongoing novelty also dovetails with the algorithmic-information view that relates strong forms of OEE to undecidability. Hernández-Orozco et al. (2018) showed that, for a broad class of sophistication-like complexity measures, decidability imposes absolute limits on stable complexity growth and that strong OEE entails undecidability/irreducibility of adapted states. While our $\Omega$ deliberately does not attempt to estimate sophistication (Kolmogorov, 1965) or logical depth (Bennett, 1988), it is consistent with that result: in regimes where undecidable structure drives unbounded production of phenotypic outcomes, $\Omega$ will continue to rise with the number of distinct attractors realized; in decidable regimes, $\Omega$ saturates. Thus, our metric provides a computable proxy for detecting the phenomenology that undecidability results lead us to expect, without invoking heavy algorithmic-complexity estimators whose incapability of distinguishing novelty from noise has been shown (Posobin and Shen, 2018).
As shown in the Supplementary Material, the external benchmarks and baseline comparisons reinforce the idea that $\Omega$ is a recurrence-weighted persistence diagnostic rather than a raw novelty or entropy score. In the ECA panel (Fig. S23), fixed-point rules yield $\Omega(T)\approx 1/T$ (rapid settling), while Rule 54 attains the largest average value, consistent with its well-known Class-IV localized-structure dynamics on finite periodic lattices. Interestingly, additive Rules 90 and 150 also score highly because finite-ring linear recurrences can generate exact long-period returns, whereas Rule 110, although also Class-IV, yields smaller $\Omega$ under our random initial condition protocol because exact whole-lattice recurrences are not reached as efficiently within the finite window. A fuller explanation of how Wolfram class, additivity, lattice size, and initial-condition ensemble jointly determine $\Omega$ lies beyond the scope of this paper.
Baseline proxy comparisons (Fig. S24) further show that $\Omega$ is not reducible to unique-state growth, entropy, or compressibility alone: high novelty/noise can coexist with low $\Omega$, whereas structured recurrence yields higher $\Omega$ even when novelty proxies are moderate. Finally, the complement-quotient control leaves the qualitative conclusions unchanged in the external benchmark analyses, addressing the concern that attractor diversity could be inflated by trivial Boolean relabelings. The curated Boolean GRN benchmark (Fig. S25) should likewise be read as a scale comparison rather than as a phase-classification exercise: for small deterministic controls, $\Omega$ mainly reflects realized finite-horizon cycle lengths, whereas the biological point is that curated GRNs remain close to the finite-horizon floor.
@@
Furthermore, our metric is squarely aimed at the York workshop’s category of “ongoing generation of adaptive novelty” (Taylor et al., 2016). It implements exactly what that category asks us to detect: sustained production of new (and potentially adaptive) phenotypes over time, operationalized in finite windows and applicable across substrates. We explicitly do not promote York’s second category (“ongoing growth of complexity”) as a necessary lens on OEE. Beyond long-standing conceptual worries about what counts as “complexity,” there is little biological support for the claim that natural selection relentlessly drives increases in organismal complexity; recent arguments review both theory and evidence indicating that complexity increases are contingent, often nonadaptive, and concentrated in a few lineages, with many trends pointing in the opposite direction (Lynch, 2025). Our focus on novelty avoids conflating open-endedness with monotone complexity growth, in line with these cautions.